Previous | Calendar |
November |
2007 |
Next | ||
---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday |
28JD301 |
29JD302Haze from CA fires present. |
30JD303Site Visit (JB, SpB, JH, NT): Details , Photos (13) rebooted isffa downloaded from all loggers BUT cnr replaced cable for the Rnet sensor at 2m on the mini tower; see details for more info tower N2 @ 350 psi at soil logger am25t mux---added 3 more TCs to the logger (now collecting data from all 25 TCs). setup the TCs near the snowtemp probes. |
31JD304Halloween |
1JD305Site Visit (JB, NT): Details worked on last snow fence plot swapped 4 flasks |
2JD306 |
3JD307 |
4JD308Daylight savings time ends (start MST). |
5JD309 |
6JD310 |
7JD311 |
8JD312Site Visit (JB, SpB, JH, NT, SL): Details , Photos (15) retired N2 cylinder CC199706 @~100 psi, started N2 cylinder CC198686 @300 psi! from 10:45-11:40 MST, tried to fix REBs amp connector (didn't work out). from 13:08-13:20 MST, did a "conduction test" with snowtemp probes---changed program to "mrs_02b.csi" which samples every 10 sec. downloaded from cnr, mrs, tcs, and soil |
9JD313 |
10JD314Site Visit (SpB, NT, SL): Details , Photos (16) Finished FC6 install on quacker2 (dell inspiron 7500) between 11:30-13:00 MST, repaired connector for 2m REBs Rnet sensor (now seems to be working) Warm and Blustery. Some light snow flurries as we departed. |
11JD315 |
12JD316 |
13JD317 |
14JD318JB has morning Dr.'s appointment |
15JD319 |
16JD320 (very windy)Site Visit (JB): Details retired N2 cylinder CC198686 @ ~100 psi, checked pressure of then retired N2 cylinder 9149741 @ ~300 psi, hauled these back to Ramaley started Jia's old pressure bomb N2 cylinder CC125555 @ ~1275 psi (leaked checked it between 12:15 and 2:30 MST) hauled fresh N2 cylinder CC193045 out to the tower as back up attached new 12 foot orange parachute cord to the tree pruner at the shed downloaded data from mrs logger (Jia's cr10), then swapped it out and re-deployed the repaired cr10wp swapped 4 flasks at the west robot, hauled out a suitcase to ship back to Utah grabbed 5 clamps for SO, brought back to Ramaley |
17JD321Site Visit (DB, NM, DG, +): Details Dave Bowling (Utah) and Noah Molotch (UCLA) were at the site setting up winter-time stuff. noah started running "noah_m.csi" in snow depth logger. |
18JD322Site Visit (DB(?), NM): Details they re-visited the site today too? |
19JD323 (very windy)Site Visit (JB, SpB, SL, NT, MW, CW): Details , Photos (23) re-deployed the repaired/upgraded propvane at 21.5m at 12:35 MST, sent "prf_01.csi" to the PROF cr23x logger. at 12:48 MST, sent "rad_03.csi" to the CNR cr23x logger. re-deployed the repaired CSM1 card reader module at the mrs logger between 13:30-14:15 MST, changed orientation of the 2-m LI-7500 downloaded from all loggers, including Noah's cr10 |
20JD324 |
21JD325Significant Snow: started on the evening of 11/20. |
22JD326CU Holiday (Turkey Day) |
23JD327CU Holiday (Day After Turkey Day) |
24JD328 |
25JD329 |
26JD330 |
27JD331mailed PAR-Lite sensor (s/n 060859) back to OSU |
28JD332 |
29JD333 (cold, windy)Site Visit (JB, JH): Details see details for map of hydra inlet (US1, US2, UN1, UN2) locations downloaded from the soil and tc loggers, swapped the pc card at the cr10_mrs logger snow depths: facing north, l=9cm, middle=10cm, r=4cm swapped 4 flasks at the west robot checked on the span and N2 cylinders at the tower (span@350 psi, N2@850 psi) brushed snow off of quacker's box, the soil, tc, noah and mrs logger boxes, and some other boxes on the cuff tower confirmed the Netgear switch is in the trailer checked the network cable connections for the tower network and the network to the world at russter2 Jia was collecting needle and phloem samples SpB Off |
30JD334 |
1JD335 |